# Optimal Rates of Convergence for Entropy Regularization in Discounted Markov Decision Processes

### v3

@@ Abstract.
We study the error introduced by entropy regularization in infinite-horizon discrete discounted Markov decision processes. We show that this error decreases exponentially in the inverse regularization strength both in a weighted KL-divergence and in value with a problem-specific exponent. This is in contrast to previously known estimates, of the order $O(\tau)$, where $\tau$ is the regularization strength. We provide a lower bound matching our upper bound up to a polynomial term thereby characterizing the exponential convergence rate for entropy regularization. Our proof relies on the observation that the solutions of entropy-regularized Markov decision processes solve a gradient flow of the unregularized reward with respect to a Riemannian metric common in natural policy gradient methods. This correspondence allows us to identify the limit of this gradient flow as the generalized maximum entropy optimal policy, thereby characterizing the implicit bias of this gradient flow which corresponds to a time-continuous version of the natural policy gradient method. We use our improved error estimates to show that for entropy-regularized natural policy gradient methods, the overall error decays exponentially in the square root of the number of iterations, improving over existing sublinear guarantees. Finally, we extend our analysis to settings beyond the entropy. In particular, we characterize the implicit bias regarding general convex potentials and their resulting generalized natural policy gradients.
@@ 1. Introduction
Entropy regularization plays an important role in reinforcement learning and is usually employed to encourage exploration thereby improving sample complexity and improving convergence of policy optimization techniques [50, 47, 49, 43, 35], where we refer to [42] for an overview of different approaches. One benefit of entropy regularization is that it leads to a tamer loss landscape [2] and indeed it corresponds to a strictly convex regularizer in the space of state-action distribution, where the reward optimization problem is equivalent to a linear program, see [42]. Employing this hidden strong convexity, vanilla and natural policy gradient methods with entropy regularization have been shown to converge exponentially fast for discrete and continuous problems, for gradient ascent and gradient flows, and for tabular methods as well as under function approximation [34, 12, 11, 40, 27]. However, the improved convergence properties of entropy-regularization in policy optimization come at the expense of an error introduced by the regularization, for which a concise characterization remains elusive. A general theory of regularized Markov decision processes was established in [19], where for a bounded regularizer and regularization strength $\tau$ a $O(\tau)$ estimate on the regularization error is provided. This guarantee covers general regularizers but neither uses the structure of Markov decision processes nor the entropic regularizer, which is common in reinforcement learning. Combining this estimate with the $O(e^{-\tau\eta k})$ convergence guarantees of entropy-regularized natural policy gradient with tabular softmax policies this yields $O(\frac{\log k}{\eta k})$ convergence of the overall error, where $k$ denotes the number of iterations and $\tau=\frac{\log k}{\eta k}$, see [12, 45]. We exploit the structure of the reward and entropy function and provide an explicit analysis of the entropy regularization error.
@@ 1.1. Contributions
The main contribution of this article is to identify the optimal exponential convergence rates of the entropy regularization error in infinite-horizon discrete discounted Markov decision processes. More precisely, we summarize our contributions as the following:
@@
  • Implicit bias. The correspondence between the Kakade gradient flow and the solutions of the entropy-regularized problems allows us to show that the gradient flows corresponding to natural policy gradients converge towards the generalized maximum entropy optimal policy thereby characterizing the implicit bias of these methods, see Theorem 3.9
  • Optimal rates for entropy regularization. In Theorem 4.2 and Theorem 4.5 we give up to polynomial terms matching exponential $O(e^{-\Delta\tau^{-1}})$ upper and lower bounds on the entropy regularization error for a problem-dependent exponent $\Delta>0$.
@@ 1.2. Related works
A general theory of regularized Markov decision processes was established in [19]. Here, $O(\tau)$ convergence was established for bounded regularizers, which was subsequently used in the overall error analysis of entropy-regularized natural policy gradients [17, 31]. Whereas the $O(\tau)$ estimate on the regularization error holds for general regularizers, it neither employs the specific structure of Markov decision processes nor the entropy function. In contrast, we show that the optimal entropy-regularized policies solve a gradient flow and use this to provide a $\tilde{O}(e^{-\Delta\tau^{-1}})$ estimate on the entropy regularization error, where $\tilde{O}$ hides a polynomial term in $\tau^{-1}$.
@@
At the core of our argument lies the observation that the optimal entropy-regularized policies solve the gradient flow of the unregularized reward with respect to the Kakade metric, which can be seen as the continuous time limit of unregularized natural policy gradient methods for tabular softmax policies. This correspondence uses the isometry between the Kakade metric and the metric induced by the conditional entropy on the state-action distributions established in [40]. This allows us to use tools from the theory of Hessian gradient flows in convex optimization [3, 37]. A similar approach has been taken recently, for the study of Fisher-Rao gradient flows of linear programs and state-action natural policy gradients [38].
@@
Natural policy gradients have been shown to converge at a $O(\frac{1}{k})$ rate [1] which was used in [28] to show assymptotic $O(e^{-ck})$ convergence for all $c<\Delta$. Our continuous time analysis uses a similar line of reasoning, but we provide an anytime analysis, a lower bound matching up to a polynomial term, and show convergence of the policies towards the generalized maximum entropy optimal policy. Further, we generalize our anytime analysis and optimal exponential rates to time natural policy gradient methods extending the existing asymptotic upper and lower bounds given in [28, 33].
@@
A continuous-time analysis of gradient flows with respect to the Kakade metric has been conducted for Markov decision processes with discrete and continuous state and action spaces in [40, 27]. Under the presence of entropy regularization with strength $\tau>0$ exponential $O(e^{-\tau t})$ convergence was established in [40, 27], which was generalized to sublinear convergence for certain decaying regularization strengths in [45]. For unregularized reward, exponential convergence was established in [40] without control over the exponent or the coefficient. For gradient flows with respect to the Fisher metric in state-action space exponential convergence $O(e^{-\delta t})$ with a problem-specific exponent was established in [38], however, a lower bound is missing there, and the exponent $\delta\leq\Delta$ is dominated by the one for Kakade gradient flows. For Riemannian metrics corresponding to $\beta$-divergences, $O(t^{\beta^{-1}})$ convergence has been established for $\beta\in(-1,0)$ [40].
@@
The algorithmic bias of policy optimization techniques has been studied before. Here, a $O(\log k)$ bound on the distance – measured in the Kakade divergence – of the policies to the maximum entropy optimal policy was established for an unregularized natural actor-critic algorithm [21]. Working in continuous time allows us to show exponential convergence towards the generalized maximum entropy optimal policy. A similar algorithmic bias showing convergence to the maximum entropy optimal policy was established for natural policy gradient methods that decrease regularization and increase step sizes during optimization [32]. In contrast to our implicit bias result, this approach considers an asymptotically vanishing but explicit regularization.
@@ 1.3. Notation
For a set $\mathscr{X}$ we denote the free vector space over $\mathscr{X}$ by $\mathbb{R}^{\mathscr{X}}\coloneqq\{f\colon\mathscr{X}\to\mathbb{R}\}$ and similarly we define >$\mathbb{R}^{\mathscr{X}}_{>0}$ and $\mathbb{R}^{\mathscr{X}}_{\geq 0}$ as all positive and non-negative real functions on $\mathscr{X}$. For a finite set $\mathscr{X}$ we denote the support of $\mu\in\mathbb{R}^{\mathscr{X}}$ by $\operatorname{supp}(\mu)\coloneqq\{x\in\mathscr{X}:\mu(x)\neq 0\}$. For a convex differentiable function $\phi\colon\Omega\to\mathbb{R}$ defined on a convex subset $\Omega\subseteq\mathbb{R}^{\mathscr{X}}$ we denote the corresponding Bregman divergence by $D_{\phi}(\mu,\nu)\coloneqq\phi(\mu)-\phi(\nu)-\nabla\phi(\nu)^{\top}(\mu-\nu)$. We denote the Shannon entropy of a vector $\mu\in\mathbb{R}_{>0}^{\mathscr{X}}$ by $H(\mu)\coloneqq-\sum_{x\in\mathscr{X}}\mu(x)\log\mu(x)$. The Bregman divergence induced by the negative Shannon entropy $-H$ is given by Kullback-Leibler or KL-divergence that we denote by $D_{\textup{KL}}(\mu,\nu)\coloneqq\sum_{x\in\mathscr{X}}\mu(x)\log\frac{\mu(x)}{\nu(x)}$. The probability simplex $\Delta_{\mathscr{X}}\coloneqq\{\mu\in\mathbb{R}_{\geq 0}^{\mathscr{X}}:\sum_{x\in\mathscr{X}}\mu(x)=1\}$ over a finite set $\mathscr{X}$ denotes the set of all probability vectors. Given a second finite set $\mathscr{Y}$, we can identify the Cartesian product $\Delta_{\mathscr{X}}^{\mathscr{Y}}=\Delta_{\mathscr{X}}\times\cdots\times\Delta_{\mathscr{X}}$ with the set of stochastic matrices or the set of Markov kernels. We refer to $\Delta_{\mathscr{X}}^{\mathscr{Y}}$ as the conditional probability polytope and for $P\in\Delta_{\mathscr{X}}^{\mathscr{Y}}$ we write $P(x|y)$ for the entries of the Markov kernel. For $\mu\in\Delta_{\mathscr{Y}}$ and a Markov kernel $P\in\Delta_{\mathscr{X}}^{\mathscr{Y}}$ we denote the corresponding joint probability measure over $\mathscr{X}\times\mathscr{Y}$ by $\mu\ast P\in\Delta_{\mathscr{X}\times\mathscr{Y}}$ given by $(\mu\ast P)(x,y)\coloneqq\mu(x)P(y|x)$. Finally, for two vectors $\mu,\nu\in\mathbb{R}^{\mathscr{X}}$ we denote the Hadamard product by $\mu\odot\nu\in\mathbb{R}^{\mathscr{X}}$ with entries $(\mu\odot\nu)(x)\coloneqq\mu(x)\nu(x)$.
@@ 2.1. Markov decision processes and entropy regularization
We consider a finite set $\mathscr{S}$ of states of some system and a finite set $\mathscr{A}$ of actions that can used to control the state $s\in\mathscr{S}$. The transition dynamics are described by a fixed Markov kernel $P\in\Delta_{\mathscr{S}}^{\mathscr{S}\times\mathscr{A}}$, where $P(s^{\prime}|s,a)$ describes the probability of transitioning from $s$ to $s^{\prime}$ under the action $a$. It is the goal in Markov decision processes and reinforcement learning to design a policy $\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}$, where the entries $\pi(a|s)$ of the stochastic matrix describe the probability of selecting an action $a$ when in state $s$. A common optimality criterion is the discounted infinite horizon reward given by
@@
$r\in\mathbb{R}^{\mathscr{S}\times\mathscr{A}}$ is the instantaneous reward vector, and $\gamma\in[0,1)$ is the discount factor. To encourage exploration it is common to regularize the reward with an entropy term which yields the entropy-regularized or KL-regularized reward
@@
The regularization can be interpreted as a convex regularization and consequently was used to show exponential convergence [34, 12, 11, 29], however, it introduces an error and leads to a new optimal policy $\pi^{\star}_{\tau}$, which might not be optimal with respect to the unregularized reward $R$. It is our goal to understand how well the entropy-regularized reward optimization problem (2.4) approximates the unregularized reward optimization problem. For this we provide an explicit analysis of the entropy regularization error $R^{\star}-R(\pi^{\star}_{\tau})$ and $\min_{\pi^{\star}\in\Pi^{\star}}D(\pi^{\star},\pi_{\tau}^{\star})$, where $R^{\star}=\max_{\pi}R(\pi)$ denotes the optimal reward, $\Pi^{\star}$ the set of optimal policies, and $D(\cdot,\cdot)$ is some notion of distance.
@@
It is well known that there are optimal policies $\pi^{\star}_{\tau}\in\Delta_{\mathscr{A}}^{\mathscr{S}}$ satisfying $V^{\star}_{\tau}=V^{\pi^{\star}_{\tau}}$ and $Q^{\star}_{\tau}=Q^{\pi^{\star}_{\tau}}$, which is known as the Bellman principle, see [19]. In the unregularized case, the optimal advantage function $A^{\star}$ satisfies $A^{\star}(s,a)\leq 0$ for all $s\in\mathscr{S},a\in\mathscr{A}$, and an action $a\in\mathscr{A}$ is optimal in a state $s\in\mathscr{S}$ if and only if $A^{\star}(s,a)=0$. For $s\in\mathscr{S}$, we denote the set of optimal actions by $A^{\star}_{s}\coloneqq\{a\in\mathscr{A}:A^{\star}(s,a)=0\}$ and the set of optimal policies is given by
@@
Note that indeed $d^{\pi}\in\Delta_{\mathscr{S}}$ by the geometric series. Sometimes, the state distributions are called state frequencies, state occupancy measures or (state) visitation distributions. We work with the following notion of distance between policies.
@@
Note that although $D_{\textup{KL}}$ is a Bregman divergence, the Kakade divergence is not, which hinders the direct use of mirror descent tools developed in the context of convex optimization [10].
@@
Bregman divergences are well-studied in convex optimization [3, 10], however, the Kakade divergence does not fall into this class. Indeed, Bregman divergences are convex in their first argument, which is not generally true for the Kakade divergence. To construct a specific example, where $D_{\operatorname{K}}(\cdot,\pi)$ is not convex, we consider the Markov decision process with two states and actions shown in Figure 1.
@@
The following sublinear estimate on the regularization error is well known and commonly applied when approximating unregularized Markov decision processes via regularization, see [19, 12]. Here, we follow the terminology common in optimization theory, where linear convergence refers to exponential convergence $O(e^{-ct})$ and sublinear to an algebraic convergence rate $O(t^{-\kappa})$ [41].

### v5

@@ Abstract.
We study the error introduced by entropy regularization in infinite-horizon discrete discounted Markov decision processes. We show that this error decreases exponentially in the inverse regularization strength, both in a weighted KL-divergence and in value with a problem-specific exponent. This is in contrast to previously known estimates, of the order $O(\tau)$, where $\tau$ is the regularization strength. We provide a lower bound that matches our upper bound up to a polynomial term, thereby characterizing the exponential convergence rate for entropy regularization. Our proof relies on the observation that the solutions of entropy-regularized Markov decision processes solve a gradient flow of the unregularized reward with respect to a Riemannian metric common in natural policy gradient methods. This correspondence allows us to identify the limit of this gradient flow as the generalized maximum entropy optimal policy, thereby characterizing the implicit bias of this gradient flow, which corresponds to a time-continuous version of the natural policy gradient method. We use our improved error estimates to show that for entropy-regularized natural policy gradient methods, the overall error decays exponentially in the square root of the number of iterations, improving over existing sublinear guarantees. Finally, we extend our analysis to settings beyond the entropy. In particular, we characterize the implicit bias regarding general convex potentials and their resulting generalized natural policy gradients.
@@ 1. Introduction
Entropy regularization plays an important role in reinforcement learning and is usually employed to encourage exploration, thereby improving sample complexity and improving convergence of policy optimization techniques [williams1991function, sutton2018reinforcement, szepesvari2022algorithms, peters2010relative, mnih2016asynchronous], where we refer to [neu2017unified] for an overview of different approaches. One benefit of entropy regularization is that it leads to a tamer loss landscape [ahmed2019understanding], and indeed it corresponds to a strictly convex regularizer in the space of state-action distribution, where the reward optimization problem is equivalent to a linear program, see [neu2017unified]. Employing this hidden strong convexity, vanilla and natural policy gradient methods with entropy regularization have been shown to converge exponentially fast for discrete and continuous problems, for gradient ascent and gradient flows, and for tabular methods as well as under function approximation [mei2020global, cen2021fast, cayci2021linear, muller2024geometry, kerimkulov2023fisher]. However, the improved convergence properties of entropy-regularization in policy optimization come at the expense of an error introduced by the regularization, for which a concise characterization remains elusive. A general theory of regularized Markov decision processes was established in [geist2019theory], where for a bounded regularizer and regularization strength $\tau$, an $O(\tau)$ estimate on the regularization error is provided. This guarantee covers general regularizers but neither uses the structure of Markov decision processes nor the entropic regularizer, which is common in reinforcement learning. Combining this estimate with the $O(e^{-\tau\eta k})$ convergence guarantees of entropy-regularized natural policy gradient with tabular softmax policies this yields $O(\frac{\log k}{\eta k})$ convergence of the overall error, where $k$ denotes the number of iterations and $\tau=\frac{\log k}{\eta k}$, see [cen2021fast, sethi2024entropy]. We exploit the structure of the reward and entropy function and provide an explicit analysis of the entropy regularization error.
@@ 1.1. Contributions
The main contribution of this article is to identify the optimal exponential convergence rates of the entropy regularization error in infinite-horizon discrete discounted Markov decision processes. More precisely, we summarize our contributions as follows:
@@
  • Implicit bias. The correspondence between the Kakade gradient flow and the solutions of the entropy-regularized problems allows us to show that the gradient flows corresponding to natural policy gradients converge towards the generalized maximum entropy optimal policy, thereby characterizing the implicit bias of these methods, see Theorem 3.9
  • Optimal rates for entropy regularization. In Theorem 4.2 and Theorem 4.5, we give up to polynomial terms matching exponential $O(e^{-\Delta\tau^{-1}})$ upper and lower bounds on the entropy regularization error for a problem-dependent exponent $\Delta>0$.
@@ 1.2. Related works
A general theory of regularized Markov decision processes was established in [geist2019theory]. Here, $O(\tau)$ convergence was established for bounded regularizers, which was subsequently used in the overall error analysis of entropy-regularized natural policy gradients [dai2018sbeed, lee2018sparse]. Whereas the $O(\tau)$ estimate on the regularization error holds for general regularizers, it neither employs the specific structure of Markov decision processes nor the entropy function. In contrast, we show that the optimal entropy-regularized policies solve a gradient flow and use this to provide a $\tilde{O}(e^{-\Delta\tau^{-1}})$ estimate on the entropy regularization error, where $\tilde{O}$ hides a polynomial term in $\tau^{-1}$.
@@
At the core of our argument lies the observation that the optimal entropy-regularized policies solve the gradient flow of the unregularized reward with respect to the Kakade metric, which can be seen as the continuous time limit of unregularized natural policy gradient methods for tabular softmax policies. This correspondence uses the isometry between the Kakade metric and the metric induced by the conditional entropy on the state-action distributions established in [muller2024geometry]. This allows us to use tools from the theory of Hessian gradient flows in convex optimization [alvarez2004hessian, mueller2023thesis]. A similar approach has been taken recently for the study of Fisher-Rao gradient flows of linear programs and state-action natural policy gradients [muller2024fisher].
@@
Natural policy gradients have been shown to converge at a $O(\frac{1}{k})$ rate [agarwal2021theory] which was used in [khodadadian2022linear] to show assymptotic $O(e^{-ck})$ convergence for all $c<\Delta$. Our continuous time analysis uses a similar line of reasoning, but we provide an anytime analysis, a lower bound matching up to a polynomial term, and show convergence of the policies towards the generalized maximum entropy optimal policy. Further, we generalize our anytime analysis and optimal exponential rates to time natural policy gradient methods, extending the existing asymptotic upper and lower bounds given in [khodadadian2022linear, liu2024elementary].
@@
A continuous-time analysis of gradient flows with respect to the Kakade metric has been conducted for Markov decision processes with discrete and continuous state and action spaces in [muller2024geometry, kerimkulov2023fisher]. Under the presence of entropy regularization with strength $\tau>0$, exponential $O(e^{-\tau t})$ convergence was established in [muller2024geometry, kerimkulov2023fisher], which was generalized to sublinear convergence for certain decaying regularization strengths in [sethi2024entropy]. For the case of unregularized reward, exponential convergence was established in [muller2024geometry] without control over the exponent or the coefficient. For gradient flows with respect to the Fisher metric in state-action space, exponential convergence $O(e^{-\delta t})$ with a problem-specific exponent was established in [muller2024fisher], however, a lower bound is missing there, and the exponent $\delta\leq\Delta$ is dominated by the one for Kakade gradient flows. For Riemannian metrics corresponding to $\beta$-divergences, $O(t^{\beta^{-1}})$ convergence has been established for $\beta\in(-1,0)$ [muller2024geometry].
@@
The algorithmic bias of policy optimization techniques has been studied before. Here, a $O(\log k)$ bound on the distance – measured in the Kakade divergence – of the policies to the maximum entropy optimal policy was established for an unregularized natural actor-critic algorithm [hu2021actor]. Working in continuous time allows us to show exponential convergence towards the generalized maximum entropy optimal policy. A similar algorithmic bias showing convergence to the maximum entropy optimal policy was established for natural policy gradient methods that decrease regularization and increase step sizes during optimization [li2023homotopic]. In contrast to our implicit bias result, this approach considers an asymptotically vanishing but explicit regularization.
@@ 1.3. Notation
For a set $\mathscr{X}$ we denote the free vector space over $\mathscr{X}$ by $\mathbb{R}^{\mathscr{X}}\coloneqq\{f\colon\mathscr{X}\to\mathbb{R}\}$ and similarly we define $\mathbb{R}^{\mathscr{X}}_{>0}$ and $\mathbb{R}^{\mathscr{X}}_{\geq 0}$ as all positive and non-negative real functions on $\mathscr{X}$. For a finite set $\mathscr{X}$ we denote the support of $\mu\in\mathbb{R}^{\mathscr{X}}$ by $\operatorname{supp}(\mu)\coloneqq\{x\in\mathscr{X}:\mu(x)\neq 0\}$. For a convex differentiable function $\phi\colon\Omega\to\mathbb{R}$ defined on a convex subset $\Omega\subseteq\mathbb{R}^{\mathscr{X}}$ we denote the corresponding Bregman divergence by $D_{\phi}(\mu,\nu)\coloneqq\phi(\mu)-\phi(\nu)-\nabla\phi(\nu)^{\top}(\mu-\nu)$. We denote the Shannon entropy of a vector $\mu\in\mathbb{R}_{>0}^{\mathscr{X}}$ by $H(\mu)\coloneqq-\sum_{x\in\mathscr{X}}\mu(x)\log\mu(x)$. The Bregman divergence induced by the negative Shannon entropy H-H is given by Kullback-Leibler or KL-divergence that we denote by $D_{\textup{KL}}(\mu,\nu)\coloneqq\sum_{x\in\mathscr{X}}\mu(x)\log\frac{\mu(x)}{\nu(x)}$. The probability simplex $\Delta_{\mathscr{X}}\coloneqq\{\mu\in\mathbb{R}_{\geq 0}^{\mathscr{X}}:\sum_{x\in\mathscr{X}}\mu(x)=1\}$ over a finite set $\mathscr{X}$ denotes the set of all probability vectors. Given a second finite set $\mathscr{Y}$, we can identify the Cartesian product $\Delta_{\mathscr{X}}^{\mathscr{Y}}=\Delta_{\mathscr{X}}\times\cdots\times\Delta_{\mathscr{X}}$ with the set of stochastic matrices or the set of Markov kernels. We refer to $\Delta_{\mathscr{X}}^{\mathscr{Y}}$ as the conditional probability polytope and for $P\in\Delta_{\mathscr{X}}^{\mathscr{Y}}$ we write $P(x|y)$ for the entries of the Markov kernel. For $\mu\in\Delta_{\mathscr{Y}}$ and a Markov kernel $P\in\Delta_{\mathscr{X}}^{\mathscr{Y}}$ we denote the corresponding joint probability measure over $\mathscr{X}\times\mathscr{Y}$ by $\mu\ast P\in\Delta_{\mathscr{X}\times\mathscr{Y}}$ given by $(\mu\ast P)(x,y)\coloneqq\mu(x)P(y|x)$. Finally, for two vectors $\mu,\nu\in\mathbb{R}^{\mathscr{X}}$ we denote the Hadamard product by $\mu\odot\nu\in\mathbb{R}^{\mathscr{X}}$ with entries $(\mu\odot\nu)(x)\coloneqq\mu(x)\nu(x)$.
@@ 2.1. Markov decision processes and entropy regularization
We consider a finite set $\mathscr{S}$ of states of some system and a finite set $\mathscr{A}$ of actions that can be used to control the state $s\in\mathscr{S}$. The transition dynamics are described by a fixed Markov kernel $P\in\Delta_{\mathscr{S}}^{\mathscr{S}\times\mathscr{A}}$, where $P(s^{\prime}|s,a)$ describes the probability of transitioning from $s$ to $s^{\prime}$ under the action $a$. It is the goal in Markov decision processes and reinforcement learning to design a policy $\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}$, where the entries $\pi(a|s)$ of the stochastic matrix describe the probability of selecting an action $a$ when in state $s$. A common optimality criterion is the discounted infinite horizon reward given by
@@
$r\in\mathbb{R}^{\mathscr{S}\times\mathscr{A}}$ is the instantaneous reward vector, and $\gamma\in[0,1)$ is the discount factor. To encourage exploration, it is common to regularize the reward with an entropy term, which yields the entropy-regularized or KL-regularized reward
@@
The regularization can be interpreted as a convex regularization and consequently was used to show exponential convergence [mei2020global, cen2021fast, cayci2021linear, lan2022policy], however, it introduces an error and leads to a new optimal policy $\pi^{\star}_{\tau}$, which might not be optimal with respect to the unregularized reward $R$. It is our goal to understand how well the entropy-regularized reward optimization problem (2.4) approximates the unregularized reward optimization problem. For this we provide an explicit analysis of the entropy regularization error $R^{\star}-R(\pi^{\star}_{\tau})$ and $\min_{\pi^{\star}\in\Pi^{\star}}D(\pi^{\star},\pi_{\tau}^{\star})$, where $R^{\star}=\max_{\pi}R(\pi)$ denotes the optimal reward, $\Pi^{\star}$ the set of optimal policies, and $D(\cdot,\cdot)$ is some notion of distance.
@@
It is well known that there are optimal policies $\pi^{\star}_{\tau}\in\Delta_{\mathscr{A}}^{\mathscr{S}}$ satisfying $V^{\star}_{\tau}=V^{\pi^{\star}_{\tau}}$ and $Q^{\star}_{\tau}=Q^{\pi^{\star}_{\tau}}$, which is known as the Bellman principle, see [geist2019theory]. In the unregularized case, the optimal advantage function $A^{\star}$ satisfies $A^{\star}(s,a)\leq 0$ for all $s\in\mathscr{S},a\in\mathscr{A}$, and an action $a\in\mathscr{A}$ is optimal in a state $s\in\mathscr{S}$ if and only if $A^{\star}(s,a)=0$. For $s\in\mathscr{S}$, we denote the set of optimal actions by $A^{\star}_{s}\coloneqq\{a\in\mathscr{A}:A^{\star}(s,a)=0\}$ and the set of optimal policies is given by
@@
Note that indeed $d^{\pi}\in\Delta_{\mathscr{S}}$ by the geometric series. Sometimes, the state distributions are called state frequencies, state occupancy measures, or (state) visitation distributions. We work with the following notion of distance between policies.
@@
Note that although $D_{\textup{KL}}$ is a Bregman divergence, the Kakade divergence is not, which hinders the direct use of mirror descent tools developed in the context of convex optimization [bubeck2015convex].
@@
Bregman divergences are well-studied in convex optimization [alvarez2004hessian, bubeck2015convex], however, the Kakade divergence does not fall into this class. Indeed, Bregman divergences are convex in their first argument, which is not generally true for the Kakade divergence. To construct a specific example, where $D_{\operatorname{K}}(\cdot,\pi)$ is not convex, we consider the Markov decision process with two states and actions shown in Figure 1.
@@
The following sublinear estimate on the regularization error is well known and commonly applied when approximating unregularized Markov decision processes via regularization, see [geist2019theory, cen2021fast]. Here, we follow the terminology common in optimization theory, where linear convergence refers to exponential convergence $O(e^{-ct})$ and sublinear to an algebraic convergence rate $O(t^{-\kappa})$ [nesterov2018lectures].